OpenStudy (anonymous):

write an equation of variation and find k if y varies inversely as x, and y=0.3 when x=15

OpenStudy (anonymous):

start with y = k/x or y = k x substitute (x,y) values

OpenStudy (anonymous):

would it be y=k/x^n or just y=k/x

OpenStudy (anonymous):

I think "equation of variation" is y = k/x here. Adding an unknown power would give you two unknowns, without enough equations.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are very welcome.